### v1

@@ Introduction
The paper is organised as follows: The baseline case and experimental data is described in Sect. Baseline geometry and flow conditions, and the generation of the high-fidelity CFD dataset is covered in Sect. Generation of high-fidelity CFD training dataset, including descriptions of the methods and their validation. The contents and structure of the dataset are then described in Sect. Dataset contents and availability before discussing conclusions and limitations of the current work in Sect. Conclusions. Extensive additional material is provided in the Supplementary Information; SI.
@@
The open-source software library OpenFOAM (v2212) was used to solve the incompressible Navier–Stokes equations via the finite volume method (FVM) [Weller1998]. The incompressible (constant density) assumption is valid for such low Mach numbers. The high Reynolds number requires a turbulence model (see Sect. Introduction), for which a scale-resolving HRLES approach is chosen as the highest fidelity level routinely applied in the automotive industry. The model of Fuchs, Mockett et al. [mockett2015two, fuchs2014anovel, fuchs2020thegrey] is used, which a variant of the Delayed Detached-Eddy Simulation (DDES) approach of Spalart et al. [Spalart2006ANew]. It uses the Spalart-Allmaras RANS model [Spalart1992] in the near-wall region and an LES formulation equivalent to the model of Nicoud et al. [Nicoud2011] in regions of separated flow. The latter accelerates the transition from RANS to LES after separation, a problem commonly referred to as the “grey area” [Mockett2014b]. The approach has been successfully applied for numerous applications and in different CFD codes [fuchs2015assessment, fuchs2017assessment, fuchs2018recent, mockett2019industrial, fuchs2019aeroacoustic, fuchs2022two]. To ensure that RANS is active throughout the entire boundary layer, the “enhanced protection” (EP) shielding formulation from Deck & Renard’s ZDES approach [Deck2020] is additionally applied. This avoids the problem of modelled stress depletion [Menter2003a] from unwanted LES-mode activity inside attached boundary layers, which can induce premature flow separation where the near-wall grid is fine. See SI for a more detailed description of the simulation approach.
@@
The dataset is provided with the CC-BY-SA license555https://creativecommons.org/licenses/by-sa/4.0/deed.en (which permits commercial use) and is available to download at no cost through Amazon S3 without the need for an AWS account. Full details are provided in the SI and the dataset README666https://caemldatasets.s3.us-east-1.amazonaws.com/drivaer/dataset/README.txt.
@@ Conclusions
Upstream CFD’s contribution to this work was conducted partially within the EXASIM (“Exascale-fähige Softwarewerkzeuge zur Strömungssimulation im industriellen Design- und Optimierungsprozess”) project (2022-2025 https://exasim-project.com) funded by the Federal Ministry of Education and Research under grant number 16ME0677 and the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor European commission can be held responsible for them.
@@ D.1 Access to dataset
The dataset is openly accessible without any additional costs and is hosted on Amazon Web Services (AWS) using the Amazon Simple Storage Service (S3). The dataset is hosted within a S3 bucket located in the us-east-1 region (North Virginia region of the United States), thus additional latency can be expected for downloads to locations away from this geographical area.
The dataset README.txt will be kept up to date for any changes to the dataset and can be found at the following URL:
The dataset itself can be downloaded via the AWS Command Line Interface (CLI) tool, which is free-of-charge. An instruction about how to install the AWS CLI tool is given here: https://docs.aws.amazon.com/cli/latest/userguide/getting-started-install.html. After installing AWS CLI, you can use the following examples to download the dataset or subsets of it, given the full dataset is approximately 22TB.
Note: If you don’t have an AWS account you will need to add –no-sign-request within your AWS command i.e aws s3 cp –no-sign-request –recursive etc…
Example 1: Download all files (~22 TB)
aws s3 cp --recursive s3://caemldatasets/drivaer/dataset .
Example 2: Only download select files (STL, images & force and moments):
@@
#!/bin/bash
# Set the S3 bucket and prefix
S3_BUCKET="caemldatasets"
S3_PREFIX="drivaer/dataset"
# Set the local directory to download the files
LOCAL_DIR="./drivaer_data"
# Create the local directory if it doesn’t exist
mkdir -p "$LOCAL_DIR"
# Loop through the run folders from 1 to 500
for i in $(seq 1 500); do
RUN_DIR="run_$i"
RUN_LOCAL_DIR="$LOCAL_DIR/$RUN_DIR"
# Create the run directory if it doesn’t exist
mkdir -p "$RUN_LOCAL_DIR"
# Download the drivaer_i.stl file
aws s3 cp "s3://$S3_BUCKET/$S3_PREFIX/$RUN_DIR/drivaer_$i.stl" \
"$RUN_LOCAL_DIR/" --only-show-errors
# Download the force_mom_i.csv file
aws s3 cp "s3://$S3_BUCKET/$S3_PREFIX/$RUN_DIR/force_mom_$i.csv" \
"$RUN_LOCAL_DIR/" --only-show-errors
aws s3 cp --recursive "s3://$S3_BUCKET/$S3_PREFIX/$RUN_DIR/images" \
"$RUN_LOCAL_DIR/images/" --only-show-errors
done
@@ D.2 Long-term hosting/maintenance plan
The data is hosted on Amazon S3 for several reasons. Firstly it provides 11 nines of durability, i.e the data has very low risk of not being available globally from a technical point of view. Secondly, the data transfer speed can be as high as 300MB/s due to the high bandwidth Amazon network (the bottleneck will likely be on the user rather than Amazon), since a poor server performance could mean extremely long download times given the size of the dataset. Thirdly, no account or credentials are required to download the data (only AWS CLI tools, described above, which are free to download and use). Finally, there are very limited number of providers that can host such large datasets ( 22TB) and make them available in such a manner. Over the coming year we will identify additional providers to mirror this dataset, to further increase the availability in the very unlikely scenario there was a problem with the Amazon S3 storage option. In addition, a dedicated website will be created for the AhmedML [ashton2024ahmed], WindsorML [ashton2024windsor] and DrivAerML datasets to help further clarify where the data is hosted and to communicate any additional mirroring sites.
@@ D.5 DOI
At present there is no specific DOI for the dataset itself, given Amazon S3 is not a resource that can easily be assigned a DOI. However it is the intention of the authors to create a DOI once a suitable mechanism is found. For the time-being, users of the dataset will be encouraged to cite this dataset paper.
@@ E.2 Distribution
How will the dataset will be distributed (e.g., tarball on website, API, GitHub)? The dataset will be free to download from Amazon S3 (without the need for an AWS account).
When will the dataset be distributed? The dataset is already available to download via Amazon S3.
@@ E.3 Maintenance
Who will be supporting/hosting/maintaining the dataset? AWS are hosting the dataset on Amazon S3.
How can the owner/curator/manager of the dataset be contacted (e.g., email address)? The owner/curator/manager of the dataset can be contacted at neashton@amazon.com (these are also provided in the dataset README and paper).
@@
If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so? We will consider this on a case by case basis and they can contact neashton@amazon.com to discuss this further.

### v2

@@ Introduction
The paper is organised as follows: The baseline case and experimental data are first described, followed by the generation of the high-fidelity CFD dataset, including descriptions of the methods and their validation. The contents and structure of the dataset are then described, before discussing conclusions and limitations of the current work. Extensive additional material is provided in the Supplementary Information; SI.
@@
The open-source software library OpenFOAM (v2212) was used to solve the incompressible Navier–Stokes equations via the finite volume method (FVM) [Weller1998]. The incompressible (constant density) assumption is valid for such low Mach numbers. The high Reynolds number requires a turbulence model for which a scale-resolving HRLES approach is chosen as the highest fidelity level routinely applied in the automotive industry. The model of Fuchs, Mockett et al. [mockett2015two, fuchs2014anovel, fuchs2020thegrey] is used, which a variant of the Delayed Detached-Eddy Simulation (DDES) approach of Spalart et al. [Spalart2006ANew]. It uses the Spalart-Allmaras RANS model [Spalart1992] in the near-wall region and an LES formulation equivalent to the model of Nicoud et al. [Nicoud2011] in regions of separated flow. The latter accelerates the transition from RANS to LES after separation, a problem commonly referred to as the “grey area” [Mockett2014b]. The approach has been successfully applied for numerous applications and in different CFD codes [fuchs2015assessment, fuchs2017assessment, fuchs2018recent, mockett2019industrial, fuchs2019aeroacoustic, fuchs2022two]. To ensure that RANS is active throughout the entire boundary layer, the “enhanced protection” (EP) shielding formulation from Deck & Renard’s ZDES approach [Deck2020] is additionally applied. This avoids the problem of modelled stress depletion [Menter2003a] from unwanted LES-mode activity inside attached boundary layers, which can induce premature flow separation where the near-wall grid is fine. See SI for a more detailed description of the simulation approach.
@@
The dataset is provided with the CC-BY-SA license555https://creativecommons.org/licenses/by-sa/4.0/deed.en (which permits commercial use) and is available to download at no cost through HuggingFace. Full details are provided in the SI and the dataset README666https://huggingface.co/datasets/neashton/drivaerml.
@@ Conclusions
Upstream CFD’s contribution to this work was conducted partially within the EXASIM (“Exascale-fähige Softwarewerkzeuge zur Strömungssimulation im industriellen Design- und Optimierungsprozess”) project (2022-2025 https://exasim-project.com) funded by the Federal Ministry of Education and Research under grant number 16ME0677 and the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor European commission can be held responsible for them. Thank you to HuggingFace for hosting this large dataset publicly.
@@ D.1 Access to dataset
The dataset is openly accessible without any additional costs and is hosted on HuggingFace.
The dataset README will be kept up to date for any changes to the dataset and can be found at the following URL:
There is also a dedicated website which is being kept up to date:
Example 1: Download all files (~30 TB)
Please note you’ll need to have git lfs installed first, then you can run the following command:
git clone git@hf.co:datasets/neashton/drivaerml
Example 2: Only download select files (STL, & force and moments):
@@ D.2 Long-term hosting/maintenance plan
The dataset is maintained on HuggingFace - one of the most widely used locations for AI datasets. This location makes it secure for the long-term and easily integrates with most ML frameworks.
@@ D.5 DOI
A DOI will be created on HuggingFace once it is past it’s preprint stage.
@@ E.2 Distribution
How will the dataset will be distributed (e.g., tarball on website, API, GitHub)? The dataset is available on HuggingFace.
When will the dataset be distributed? The dataset is available on HuggingFace.
@@ E.3 Maintenance
Who will be supporting/hosting/maintaining the dataset? The dataset is hosted on HuggingFace and maintained by the authors (primarily Neil Ashton).
How can the owner/curator/manager of the dataset be contacted (e.g., email address)? The owner/curator/manager of the dataset can be contacted at contact@caemldatasets.org (these are also provided in the dataset README and paper).
@@
If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so? We will consider this on a case by case basis and they can contact contact@caemldatasets.org to discuss this further.